In which of the following is the oxidation state of nitrogen given incorrectly?

a. HNO(+3)
b. N2O3 (+3)
c. NaN3 (-1)
d. H2N2O2 (+1)
e. N2H4 (-2)
Answer #1
a)HNO
since oxidation state of H is1 and oxygen is -2

net oxidation state of compound to be zero
1+(x)-2=0
x=1
hence oxidation state of N is 1
b)N2O3
b)since oxidation state of oxygen is -2

net oxidation state of compound to be zero
2x+(-2)*3=0
x=3
hence oxidation state of N is 3
c)NaN3
since oxidation state of Na is +1
1+3x=0
x=-1/3
hence oxidation state of N is -1/3
d)H2N2O2

since oxidation state of H is1 and oxygen is -2
2*1+2x+(-2)*2=0
2+2x-4=0
x=1
hence oxidation state of N is 1
e)N2H4
since oxidation state of H is 1
2x+4*1=0
x=-2
hence oxidation state of N is -2
oxidation states of N that mention incorrectly are a,c